OpenStudy (anonymous):

what is the axis of symmetry of the graph of the function y=7(x+1)(x-3)?

OpenStudy (anonymous):

take the average of the zeros.... what are the zeros in your equation?

OpenStudy (anonymous):

1 and -3?

OpenStudy (anonymous):

close..

OpenStudy (anonymous):

is it the opposites?

OpenStudy (anonymous):

yes... zeros are the x-coordinates when y=0

OpenStudy (anonymous):

so your equation: y=7(x+1)(x-3) when y=0: 0=7(x+1)(x-3) solve for x..

OpenStudy (anonymous):

so -1 and 3

OpenStudy (anonymous):

once you find them, take the average to get the line of symmetry.

OpenStudy (anonymous):

yes...

OpenStudy (anonymous):

so its x=1??

OpenStudy (anonymous):

Would be worthwhile to simplify and graph this function to see how these points are related.

OpenStudy (anonymous):

so its wrong?

OpenStudy (anonymous):

@dpaInc

OpenStudy (anonymous):

nothing's wrong... what @CliffSedge is suggesting is so you can "see" where the zeros are in relation to the axis of symmetry... i suggest that too...

OpenStudy (anonymous):

oh ok, I understand. thanks both for the help (:

OpenStudy (anonymous):

then u can see why you take the average of the zeros...

OpenStudy (anonymous):

Yeah, 'cause I know I'm a visual guy. All the analytical stuff is great, but if I can *see* it, then it makes a lot more sense.

OpenStudy (anonymous):

same here... visual..

OpenStudy (anonymous):

yeah I graphed it and I see what u guys mean,thanks again